OpenStudy (anonymous):

Diane runs 10 miles in 75 minutes. At the same rate, how many miles would she run in 72 minutes?

OpenStudy (anonymous):

10 miles / 75 minutes = x miles / 72 minutes 10/75 = x/72 72*10/75 = x 9.6 miles = x